Graphing using slope-intercept form
1. x=2y-6=0
2. -x+y=-5
3. -3x+y=-2
4. -2x+y=0
5. x=2y-6
6. x+y=-5
7. 6x+2y=-4
8. 2y=-4x

Answer by Alan3354(69443)

Graphing using slope-intercept form
1. x=2y-6=0 Too many equal signs.
-------------------------------------
2. -x+y=-5
Put in slope-intercept form y = mx + b
That means solve for y
y = x - 5
Slope m = 1
Y-int = -5
---
Plot the point (0,-5)
Move 1 square right and 1 square up, plot another point.
Draw a line thru them.